# A metrological framework for uncertainty evaluation in machine learning classification models

### v2

@@ 1.3 Summary of contributions
  • We propose a metrological conceptual uncertainty evaluation framework for ML classification inspired by existing vocabularies and frameworks from the fields of qualitative chemical analysis [bettencourt_da_silva_eurachemcitac_2021] and clinical laboratory sciences [nordin_vocabulary_2018]. It is based upon the determination of the PMF of the nominal variable returned by the ML classifier, and the expression of uncertainty in terms of summary statistics derived from it. We review some possible summary statistics, discussing their benefits and drawbacks. The framework represents a natural analogue to the determination of the PDF for a quantitative variable and the derivation from it of summary statistics such as standard uncertainty. We also show how conditional probability can be used to enable uncertainty evaluation for measurement models that have categorical variables as inputs.
@@
We present our conceptual framework in Sections 2 and 3. In Section 2, we show how the ML classification task can be viewed as the realisation of a nominal variable (nominal property), and we describe how such properties are addressed in metrology. In Section 3, we present our framework for uncertainty evaluation of nominal properties based on discrete probability distributions and summary statistics. We also address the propagation of uncertainty in ongoing traceability chains. We present our two case studies illustrating the framework in Section 4, before concluding in Section 5.
2 Examination of nominal properties
2.1 Nominal variables
Classification tasks deal with nominal variables. To explain what nominal variables are, we will use Stevens’ theory of scales of measurement [stevens_theory_1946], which is frequently cited to describe the nature of different types of variables involved in measurement. We note at the outset that the question as to whether the term measurement is applicable to all types of variables is a subject of extensive debate in the metrology community and return to this question later in the section. For a deeper review of Stevens’ theory and its relevance to metrology refer to [white_meaning_2011]; see also other references which propose alternative scales of measurement to Stevens’ theory, e.g., [Mosteller1977], not considered in our study.
Stevens’ theory distinguishes four scales of measurement: nominal, ordinal, interval and ratio (see Table 1). A nominal scale has multiple categories (or classes) with no natural ordering, for example a biological species or a blood group. Although these categories can be represented by numbers, the numbers do not have any quantitative meaning. Variables measured on a nominal scale can only be compared based on their equality, i.e., same or different.
An ordinal scale also has multiple categories, but with a natural ordering. For example, wind force on the Beaufort scale or stages of cancer. Variables measured on an ordinal scale can be compared based on their equality and their order, e.g., higher/lower, stronger/weaker. Distances between categories, however, are not necessarily meaningful. For example, the difference in the severity of cancer stages I and II is not the same as the difference in the severity of cancer stages III and IV [ISO33406].
Unlike for an ordinal scale, for an interval scale the differences are meaningful. Examples of variables measured on an interval scale include longitude and latitude, year, etc. A distinctive feature of this scale is that it does not have a meaningful (‘true’) zero, which is chosen arbitrarily.
Variables measured on a ratio scale do have a ‘true’ zero and represent the most encountered variables in physics, for example length, mass, and electrical power.
@@ 2.1 Nominal variables
One of the important aspects of the scales of measurement proposed by Stevens is that different scales have different statistics representing central tendency and dispersion. From a statistical point of view, a nominal scale represents the lowest (the most primitive) scale, while a ratio scale represents the highest scale. Statistical measures that apply to the lower scales also apply to the higher scales, but not the other way around [Agresti2002]. For example, the central tendency of variables measured on a nominal scale can only be expressed by the mode, whereas the central tendency of variables measured on an ordinal scale can be expressed by the mode and the median. The notions of central tendency and dispersion are important for understanding how to express uncertainties. In Section 3, we discuss expressions of uncertainty for nominal variables.
The nomenclature for nominal variables varies across scientific communities. In the mathematical sciences, nominal and ordinal variables are often merged into a single group, categorical variables [Agresti2002, Hastie2008], with nominal variables sometimes referred to as unordered categorical variables and ordinal variables referred to as ordered categorical variables [Hastie2008]. In addition, nominal variables are sometimes called qualitative variables, and are distinguished from quantitative variables (variables measured on interval and ratio scales), with some ambiguity around the classification of ordinal variables [Agresti2002]. However, in some international standards, the term qualitative variable (or qualitative property) is used as an umbrella term for both nominal and ordinal variables [ISO33406].
In this paper, we use the term categorical variable as an umbrella term for nominal and ordinal variables. We do not use the term qualitative variable to avoid confusion because, as mentioned above, there is disagreement about whether ordinal variables are included [Agresti2002, ISO33406]. We also use interchangeably the term variable (more common in the language of data science) and the term property (more common in the language of metrology).
Below, we will discuss how the evaluation of categorical (more specifically, nominal) variables is currently addressed in
  • Authoritative metrological documents, namely the VIM and the GUM suite of documents;
  • International standards;
  • Metrology-related publications.
2.2 How nominal variables are addressed in…
2.2.1 Authoritative metrological documents
@@
property of a phenomenon, body, or substance, where the property has no magnitude
EXAMPLE 1 Sex of a human being.
EXAMPLE 2 Colour of a paint sample.
EXAMPLE 3 Colour of a spot test in chemistry.
EXAMPLE 4 ISO two-letter country code.
EXAMPLE 5 Sequence of amino acids in a polypeptide.
NOTE 1 A nominal property has a value, which can be expressed in words, by alphanumerical codes, or by other means.
NOTE 2 ‘Nominal property value’ is not to be confused with nominal quantity value. [^1] In the VIM3, a nominal quantity value is defined as “the rounded or approximate value of a characterizing quantity of a measuring instrument or measuring system that provides guidance for its appropriate use”.
Interestingly, in the VIM3, ordinal variable is considered to represent a type of quantity. Correspondingly, it is referred to as an ordinal quantity and is defined as a [JCGMVIM3, Definition 1.26]
quantity, defined by a conventional measurement procedure, for which a total ordering relation can be established, according to magnitude, with other quantities of the same kind, but for which no algebraic operations among those quantities exist
EXAMPLE 1 Rockwell C hardness.
EXAMPLE 2 Octane number for petroleum fuel.
EXAMPLE 3 Earthquake strength on the Richter scale.
EXAMPLE 4 Subjective level of abdominal pain on a scale from zero to five.
NOTE 1 Ordinal quantities can enter into empirical relations only and have neither measurement units nor quantity dimensions. Differences and ratios of ordinal quantities have no physical meaning.
NOTE 2 Ordinal quantities are arranged according to ordinal quantity-value scales (see 1.28).
In addition to providing definitions for nominal properties and ordinal quantities, the VIM3 clarifies that the term measurement relates only to quantities. In the VIM3, measurement is defined as [JCGMVIM3, Definition 2.1]the process of experimentally obtaining one or more quantity values that can reasonably be attributed to a quantity” with NOTE 1 explicitly stating that “measurement does not apply to nominal properties”. The VIM3 does not provide a separate entry for a term representing a measurement-like analogue for nominal properties. However, in Note 1 supplementing the definition of reference material [JCGMVIM3, Definition 5.13], it says “Examination of a nominal property provides a nominal property value and associated uncertainty. This uncertainty is not a measurement uncertainty”, which implies that such an analogue is examination.
In the preparation of ${}^{\text{th}}$ edition of the VIM (the VIM4), the Joint Committee for Guides in Metrology (JCGM) is holding a discussion on including an entire chapter on nominal properties, motivated by the growing interest in a metrologically rigorous handling of these properties. These properties are indeed quite common in chemistry, forensics and clinical laboratory testing. An initial draft of the chapter is largely based on the vocabulary developed for clinical laboratory science by the IFCC (International Federation of Clinical Chemistry and Laboratory Medicine) and IUPAC (International Union of Pure and Applied Chemistry) [nordin_vocabulary_2018].
Similarly to the VIM3, the vocabulary developed by the IFCC-IUPAC [nordin_vocabulary_2018] differentiates measurement from examination. Unlike the VIM3, it provides an explicit definition of examination as “the process of experimentally obtaining one or more values that can reasonably be attributed to a nominal property together with any other available relevant information”, which is drafted analogously to the VIM3 definition of measurement. The IFCC-IUPAC vocabulary also introduces a term examinand (“nominal property intended to be examined”) analogous to the VIM3 definition of measurand (“quantity intended to be measured”). It is worth noting that in some other metrology-related publications, an additional term –– evaluation –– has been used as an umbrella term, which covers both measurement and examination [mari_foundations_2020], and the term evaluand used correspondingly to represent a property to be evaluated [mari_toward_2017].
Whether the JCGM will adopt the terminology on nominal properties from the IFCC-IUPAC vocabulary [nordin_vocabulary_2018] in the published version of the VIM4 is yet to be seen. However, the metrology community seems to be divided regarding whether two separate terms –– measurement and examination –– are necessary, or whether evaluation of nominal properties should be viewed as a type of measurement. In this latter direction, some publications by the national metrology institute of the United States of America, NIST (National Institute of Standards and Technology), use the term measurement and the term measurand for both quantitative and qualitative properties [possolo_brief_2024]. Similarly, in [white_meaning_2011] –– a publication by the Measurement Standards Laboratory of New Zealand — a similar position is adopted, that the definition of measurement should be more inclusive and should cover evaluation of nominal variables in addition to quantities.
To summarise, the VIM3 provides explicit definitions of nominal properties and ordinal quantities. It also mentions that the term measurement does not apply to nominal properties, and the term examination should be used instead. However, it does not provide a separate entry for the definition of examination. The next edition of the VIM (VIM4) might bring more clarity as there is an ongoing discussion on including an entire chapter on the terminology related to nominal properties.
Unlike the VIM, the GUM suite of documents is entirely focused on quantities, with the only exception being the fourth document of the GUM suite dedicated to conformity assessment [JCGM106-2012]. As previously stated in Section 1, in conformity assessment a measured quantity is converted into a nominal variable (either conforming and non-conforming). The relationship of conformity assessment to evaluation of nominal properties was previously discussed in guidelines produced by EURACHEM (European network of organizations focusing on analytical chemistry and quality related issues) and CITAC (The Cooperation on International Traceability in Analytical Chemistry) [bettencourt_da_silva_eurachemcitac_2021].
2.2.2 International standards
@@
Although the VIM and the GUM currently provide negligible coverage of nominal variables, these variables have been addressed in a number of standards and recommendations produced by international bodies working in various areas of science and industry, where nominal variables are frequently used. As described above, a vocabulary dedicated to nominal variables – ”Vocabulary on nominal property, examination, and related concepts for clinical laboratory sciences (IFCC-IUPAC Recommendations 2017)” [nordin_vocabulary_2018] – was developed by the International Federation of Clinical Chemistry and Laboratory Medicine (IFCC) and International Union of Pure and Applied Chemistry (IUPAC). Also, nominal variables and approaches to the evaluation of uncertainties associated with such variables were discussed in ”Assessment of performance and uncertainty in qualitative chemical analysis (EURACHEM/CITAC Guide)” [bettencourt_da_silva_eurachemcitac_2021]. Recommendations on handling nominal variables specifically related to reference materials, which have nominal properties, were recently systemised in the International Standard ISO 33406:2024 ”Approaches for the production of reference materials with qualitative properties” [ISO33406]. This ISO standard provides a discussion on establishing metrological traceability for reference material with nominal properties and evaluation of their uncertainties.
In relation to traceability of reference materials with nominal properties, ISO 33406:2024 advises that well-established guidelines on metrological traceability provided for quantities (e.g., ISO/IEC 17025 ”Testing and calibration laboratories” [ISO17025]) are followed whenever it is possible. Specifically, the metrological traceability guidelines should be used when the assignment of a nominal property relies on a measurement of a quantity (e.g., melting point or pH) and/or is performed in conditions that can be quantitatively assessed (e.g., times or temperatures) If, however, the assignment of a nominal property relies on the comparison with reference data or another reference material, which is already value-assigned for the nominal property in question, then this reference should be stated in the relevant documentation accompanying the reference material.
2.2.3 Metrology-related publications
The metrology community, including national metrology institutes and members of the JCGM working groups, has also been bringing more attention to nominal variables in its publications over the past years. Examples of the use of nominal variables in metrology and discussion on their uncertainty evaluation can be found in NIST reports (e.g. [possolo_simple_2015]). Evaluation of uncertainty for nominal variables was also previously discussed by members of JCGM Working Group 2 in [mari_foundations_2020]. It is important to note that although many of the above references address evaluation and expression of uncertainties associated with nominal variables, no agreement has yet been reached on this topic. And, to our knowledge, the discussion on the definition of uncertainty associated with nominal variables to be included in the VIM4 is still ongoing.
2.2.4 Summary
In recent years, interest in nominal properties has been growing in the metrology community, as reflected in several publications, standards and most importantly in the current work on the next edition of the VIM. However, no agreement has yet been reached on whether the term ‘measurement’ should cover nominal properties and on the expression of uncertainties associated with nominal variables. Section 3 will specifically address the latter.
@@ 3 Uncertainty evaluation for nominal properties
In this section, we outline approaches for uncertainty evaluation for nominal properties within ML classification models and further propagation through measurement models. In Section 3.1, we define PDFs and PMFs, illustrating that PMFs are the natural probability distribution for nominal properties. In Section 3.2, we describe various statistics that could be used to express uncertainty of nominal variables, highlighting their strengths and weaknesses. In Section 3.3, we show, with use of the PMF, how uncertainties of nominal properties can be propagated through measurement models having categorical inputs.
@@ 3.1 Probability distributions
As stated in the GUM [JCGMGUM] and other metrology-related publications [possolo_simple_2015], the uncertainty associated with any variable is fully expressed by its probability distribution, which is the distribution around a measured value (often taken as the mean of the distribution) in which the ‘true’ value is expected to lie. This distribution can either be empirical (i.e., based upon the frequency of occurrence of the variable), or have an assumed form (such as Gaussian). For continuous quantitative variables $\mathbf{x}$ that can take values in the set $\mathcal{X}\subset\mathbb{R}^{D}$, where $D$ is the dimension of the continuous input space, the probability distribution is defined in terms of a probability distribution function (PDF), which is a map or function
@@
Condition (3.2) ensures that $p(\mathbf{x})$ is properly normalised and can act as a PDF.
Similarly, for a nominal variable with $K$ possible classes (or categories), that can take values in the set $\mathcal{C}_{K}=\{c_{1},\dots,c_{K}\}$, the probability distribution is defined in terms of the PMF:
@@
Condition (3.4) ensures $p$ obeys the correct probability sum rule and can act as a PMF.
@@
Given the analogy with PDF for quantitative variables, the PMF is the natural candidate for expressing uncertainty in nominal properties and has been motivated in metrology-related publications such as [mari_foundations_2020, possolo_brief_2024], as well as the recent international standard [ISO33406] introduced in Section 2.2.2.
@@ 3.2 Expressions of uncertainty
Expressions of uncertainty for quantitative variables such as the standard deviation are well known in the metrology community, and are used in the GUM [jcgm-2023gum1]. The standard deviation is one example of a summary statistic which expresses the uncertainty of a quantitative variable in terms of a single positive real number. More formally, given a quantitative variable $\mathbf{X}$ with PDF $p(\mathbf{x})$, a statistic $u$ for uncertainty is evaluated using a function $F$ such that
$u=F(p(\mathbf{x})),\quad u\in[0,\infty).$ (3.6)
Statistics calculated from a function $F$, such as the standard deviation or interquartile range (IQR), are measures of dispersion about a central tendency (see Table 1). Defining a statistic of dispersion for nominal properties is more challenging as the only statistic of central tendency is the mode, for which there is no natural notion of distance [mari_foundations_2020]. Thus, the more common way to express uncertainty of nominal properties is in terms of statistics (or indices) of variation [wilcox_indices_1973, mari_foundations_2020]. However, there is debate in the metrology community about the nomenclature for describing various uncertainty statistics of nominal properties in terms of variation or dispersion, such as the information entropy [nordin_vocabulary_2018, mari_foundations_2020] (see Section 3.2.2 for a definition). In this section we will avoid use of such terminology, and focus on the mathematical properties and their implications as candidates for expressions of uncertainty in nominal properties.

### v3

@@ 1.3 Summary of contributions
  • We propose a metrological conceptual uncertainty evaluation framework for nominal properties which is applicable to ML classification. Our framework is inspired by existing vocabularies and frameworks from the fields of qualitative chemical analysis [28] and clinical laboratory sciences [29]. It is based upon the determination of the PMF of the nominal variable returned by the ML classifier, and the expression of uncertainty in terms of summary statistics derived from it. We review some possible summary statistics, discussing their benefits and drawbacks. The framework represents a natural analogue to the determination of the PDF for a quantitative variable and the derivation from it of summary statistics such as standard uncertainty. We also show how conditional probability can be used to enable uncertainty evaluation for measurement models that have quantitative and categorical variables as inputs.
@@
1.4 Outline of contents
In Section 2, we give background on nominal properties, and we describe how such properties are addressed in metrology. In Section 3, we present our framework for uncertainty evaluation of nominal properties based on discrete probability distributions and summary statistics. We also address the propagation of uncertainty through multistage measurement models. We present our two ML case studies illustrating the framework in Section 4, before concluding in Section 5.
2 Nominal variables
2.1 Nominal versus other variables
Classification tasks deal with nominal variables. To explain what nominal variables are, we will use Stevens’ theory of scales of measurement [30]. For a deeper review of Stevens’ theory and its relevance to metrology we refer the reader to [31]. Stevens’ theory distinguishes four types of scales of measurement: nominal, ordinal, interval and ratio (see Table 1). A nominal scale has multiple categories (or classes) with no natural ordering, for example a biological species or a blood group. Although these categories can be represented by numbers, the numbers do not have any quantitative meaning. Properties measured on a nominal scale can only be compared based on their equivalence, i.e., same or different.
An ordinal scale also has multiple categories, but with a natural ordering, for example wind force on the Beaufort scale or stages of cancer. Properties measured on an ordinal scale can be compared based on their equality and their order, e.g., higher/lower, stronger/weaker. Distances between categories, however, are not necessarily meaningful. For example, the difference in the severity of cancer stages I and II is not the same as the difference in the severity of cancer stages III and IV [32].
Unlike for an ordinal scale, for an interval scale the differences are meaningful. Examples of properties measured on an interval scale include longitude and latitude, year, etc. A distinctive feature of this scale is that it does not have a meaningful (‘true’) zero, which is chosen arbitrarily.
Properties measured on a ratio scale do have a ‘true’ zero and represent the most encountered properties in physics, for example length, mass, and electrical power.
@@ 2.1 Nominal versus other variables
One of the important aspects of Stevens’ theory is that different types of scales have different statistics representing central tendency and dispersion. From a statistical point of view, a nominal scale represents the lowest (the most primitive) scale, while a ratio scale represents the highest scale. Statistics that apply to the lower scales also apply to the higher scales, but not the other way around [33]. The central tendency of properties measured on a nominal scale can only be expressed by the mode (although strictly speaking, the mode is representative of the most frequent class, not the central class). The central tendency of properties measured on an ordinal scale can be expressed by the mode and the median. The notions of central tendency and dispersion are important for understanding how to express uncertainties. In Section 3, we discuss expressions of uncertainty for nominal variables.
The nomenclature for nominal variables varies across scientific communities. In the mathematical sciences, nominal and ordinal variables are often merged into a single group, categorical variables [33, 20], with nominal variables sometimes referred to as unordered categorical variables and ordinal variables referred to as ordered categorical variables [20]. In addition, nominal variables are sometimes called qualitative variables, and are distinguished from quantitative variables (variables measured on interval and ratio scales), with some ambiguity around the classification of ordinal variables [33]. However, in some international standards, the term qualitative variable (or qualitative property) is used as an umbrella term for both nominal and ordinal variables [32].
In this paper, we use the term categorical variable as an umbrella term for nominal and ordinal variables. We do not use the term qualitative variable to avoid confusion because, as mentioned above, there is disagreement about whether ordinal variables are included [33, 32]. We also use interchangeably the term variable (more common in the language of data science) and the term property (more common in the language of metrology).
2.2 Nominal variables in metrology
@@
Unlike the VIM, the GUM suite of documents is entirely focused on quantities, with the only exception being the fourth document of the GUM suite dedicated to conformity assessment [25], previously mentioned in Section 1.
@@
Although the VIM and the GUM currently provide negligible coverage of nominal variables, these variables have been addressed in a number of standards and recommendations produced by international bodies working in various areas of science and industry, where nominal variables are frequently used. For example, the International Federation of Clinical Chemistry and Laboratory Medicine (IFCC) and International Union of Pure and Applied Chemistry (IUPAC) developed “Vocabulary on nominal property, examination, and related concepts for clinical laboratory sciences (IFCC-IUPAC Recommendations 2017)” [29]. Also, nominal variables and approaches to the evaluation of uncertainties associated with such variables were discussed in “Assessment of performance and uncertainty in qualitative chemical analysis (EURACHEM/CITAC Guide)” [28]. Recommendations on handling nominal variables specifically related to reference materials, which have nominal properties, were recently systemised in the International Standard ISO 33406:2024 “Approaches for the production of reference materials with qualitative properties” [32]. Discussion on nominal variables and their uncertainty evaluation can also be found in publications by national metrology institutes (e.g., [34]), as well as in publications by the members of JCGM Working Group 2  [35]. It is important to note that although many of the above references address evaluation and expression of uncertainties associated with nominal variables, no agreement has yet been reached on this topic.
@@ 3 Uncertainty evaluation for nominal properties
In this section, we outline approaches for uncertainty evaluation of nominal properties and further propagation through measurement models. In Section 3.1, we define PDFs and PMFs, illustrating that PMFs are the natural probability distribution for nominal properties. In Section 3.2, we describe various statistics that could be used to express uncertainty of nominal variables, highlighting their strengths and weaknesses and the extent of their departure from accepted notions of uncertainty in the metrology community. In Section 3.3, we show, with use of the PMF, how uncertainties of nominal properties can be propagated through measurement models having categorical inputs.
@@ 3.1 Probability distributions
As stated in the GUM [10] and other metrology-related publications [34], the uncertainty associated with any variable is fully expressed by its probability distribution, which is the distribution around a measured value (often taken as the mean of the distribution) in which the ‘true’ value is expected to lie. This distribution can either be empirical (i.e., based upon the frequency of occurrence of the variable), or have an assumed form (such as Gaussian). For continuous quantitative variables $\mathbf{x}$ that can take values in the set $\mathcal{X}\subset\mathbb{R}^{D}$, where $D$ is the dimension of the continuous input space, the probability distribution is defined in terms of a PDF [36][37, Definition 2.26], which is a map or function
@@
Condition (3.2) ensures that expression (3.1) is properly normalised and can act as a PDF.
Similarly, a nominal variable can take values in the set $\mathcal{C}_{K}=\{c_{1},\dots,c_{K}\}$, where $K$ is the number of possible classes (or categories). Thus, the probability distribution is defined in terms of the PMF [16][37, Definition 2.24]:
@@
Condition (3.4) ensures the expression (3.1) obeys the correct probability sum rule and can act as a PMF.
@@
Given the analogy with PDF for quantitative variables, the PMF is the natural candidate for expressing uncertainty in nominal properties and has been motivated in metrology-related publications such as [35, 38], as well as the recent international standard [32] introduced in Section $2.$
@@ 3.2 Expressions of uncertainty
Expressions of uncertainty for quantitative variables such as the standard deviation are well known in the metrology community, and are used in the GUM [10]. The standard deviation is one example of a summary statistic that expresses the uncertainty of a quantitative variable in terms of a single positive real number. More formally, given a quantitative variable $\mathbf{X}$ with PDF $\mathbf{p}(\mathbf{x})$, a statistic $u$ for uncertainty is evaluated using a function $F$ such that
$u=F(\mathbf{p}(\mathbf{x})),\quad u\in[0,\infty).$ (3.6)
The definition of uncertainty according to the VIM [9, 2.26] is, a “non-negative parameter characterizing the dispersion of the quantity values being attributed to a measurand, based on the information used.” A statistic of dispersion is typically assumed to involve some notion of distance, and it is has been questioned whether such a notion exists for nominal properties [35]. The PMF of a nominal property can be interpreted in terms of expected distance with respect to a simple metric. Using the Hamming distance on a single categorical variable [16, Chapter 25]
$s_{jk}=\begin{cases}0&c_{j}=c_{k},\\ 1&\text{otherwise},\end{cases}\quad j,k\in\{1,\dots,K\},$ (3.7)
the expected distance to class $c_{k}$ of a nominal property following the distribution given by $\mathbf{p}$ is given by
$\sum_{j=1}^{K}p_{j}s_{jk}=\sum_{j\neq k}p_{j}=1-p_{k}.$ (3.8)
Various so-called statistics of variation [39] have been proposed for nominal properties, and we present an overview of some of these in Section 3.2.3. Given our observation in the previous paragraph that all summary statistics of a PMF have a distance interpretation, we will refer to them throughout as statistics of dispersion.
The VIM also defines a notion of standard measurement uncertainty for quantities, namely the standard deviation, which captures dispersion about a central tendency, namely the mean [9, (2.30)]. For nominal properties, the only statistic of central tendency is the mode. Similarly to the notion of distance in (3.7), we can also define a meaningful notion of distance from the modal class (or classes in the multimodal case) for nominal properties (see Appendix A). Equipped with this notion of distance, some of the statistics we present in Section 3.2.3 can indeed be viewed as statistics of dispersion about the modal class(es). Writing $\widehat{p}=\max(\mathbf{p})$ for the probability assigned to the modal class(es), also known as the modal probability, those statistics that can be expressed as a function of $\widehat{p}$ alone can be interpreted as dispersion about central tendency.
That said, some of the statistics we present in Section 3.2.3 are not in general statistics of dispersion from a central tendency. Such statistics do not, therefore, represent a counterpart to the notion of standard measurement uncertainty for quantities.
Such a distinction is only relevant for multi-class nominal variables ($K>2$). For binary nominal variables ($K=2$), all statistics of dispersion (from the mode or otherwise) can be expressed as a function of $\widehat{p}$ alone, and so capture dispersion about central tendency. For each of the statistics of dispersion defined in Section 3.2.3, we state their explicit reduction to a function of $\widehat{p}$ alone in the binary case.
3.2.1 Uncertainty and confidence